### v2

@@ Abstract
We explore the relationship between two popular concepts in the allocation of divisible items: competitive equilibrium (CE) and allocations that maximize Nash welfare, i.e., allocations where the weighted geometric mean of the utilities is maximal. When agents have homogeneous concave utility functions, these two concepts coincide: the classical Eisenberg–Gale convex program that maximizes Nash welfare over feasible allocations yields a competitive equilibrium. However, these two concepts diverge for non–homogeneous utilities. From a computational perspective, maximizing Nash welfare amounts to solving a convex program for any concave utility functions, whereas computing CE becomes PPAD-hard already for separable piecewise linear concave (SPLC) utilities.
We introduce the concept of Gale-substitute utility functions, which is an analogue of the weak gross substitutes (WGS) property for the so-called Gale demand system. For Gale-substitutes utilities, we show that any allocation maximizing Nash welfare provides an approximate-CE with surprisingly strong guarantees, where every agent gets at least half the maximum utility they can get at any CE, and is approximately envy-free. Gale-substitutes include utility functions where computing CE is PPAD hard, such as all separable concave utilities and the previously studied non-separable class of Leontief-free utilities. We introduce a broad new class of utility functions called generalized network utilities based on the generalized flow model. This class includes SPLC and Leontief-free utilities, and we show that all such utilities are Gale-substitutes.
Conversely, although some agents may get much higher utility at a Nash welfare maximizing allocation than at a CE, we show a price of anarchy type result: for general concave utilities, every CE achieves at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight.
@@ 1 Introduction
Competitive (market) equilibrium is a central solution concept in economics to analyze and predict the outcomes of market interactions. The Fisher model is a fundamental market model where a set of $n$ agents ${\mathcal{A}}$ arrive at the market with a fixed supply of divisible goods ${\mathcal{M}}$. Each agent $i$ has a fixed budget $b_{i}$ and preferences over bundles of goods given by a utility function $u_{i}(.)$. At competitive equilibrium (CE), prices $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ of goods and fractional allocation $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ of goods to agents are such that each agent $i$ gets an optimal bundle of goods, i.e., $x_{i}$ maximizes $u_{i}(x_{i})$ subject to $\langle\mathbf{p},x_{i}\rangle\leq b_{i}$, and the market clears, i.e., the demand for each good meets its supply.
Equilibrium allocations satisfy many desirable properties, including Pareto optimality and the fairness guarantee of (weighted) envy-freeness. These properties make this an attractive concept also in other settings where no direct trade and money are involved. For example, competitive equilibrium with equal incomes (CEEI) is a popular mechanism in fair division, where a fictitious market is created by giving each agent the same budget Varian (1974). We note that CEEI is a special case of the Fisher model, where $b_{i}=1$ for each agent $i$.
It is standard to assume that the utility functions $u_{i}(.)$ are monotone non-decreasing, concave, and satisfy $u_{i}(\mathbf{0})=0$. Classical results in mathematical economics established the existence of an equilibrium in the Fisher model as well as more general exchange market models, by Arrow and Debreu (1954) and by McKenzie (1954). These results are based on fixed point theorems and thus do not lead to efficient computational methods. Finding an equilibrium by efficient computation has been an important topic in economics, optimization, and computer science over the past decades.
Eisenberg and Gale (1959) gave a convex programming formulation for CE with linear utility functions $u_{i}(x)=\sum_{j\in{\mathcal{M}}}v_{ij}x_{j}$. This program amounts to maximizing the weighted Nash welfare, i.e., the weighted geometric mean of agents’ utilities over feasible allocations. That is, any allocation maximizing $\prod_{i}u_{i}(x_{i})^{b_{i}}$ (equivalently, $\sum_{i}b_{i}\log{u_{i}(x_{i})}$) over feasible allocations gives a CE, where the dual solution yields the corresponding prices $\mathbf{p}$. Eisenberg (1961) showed that the same convex program yields competitive equilibria more generally for concave utility functions that are positively homogeneous of degree one, i.e., $u_{i}(\alpha x_{i})=\alpha u_{i}(x_{i})$ for any $\alpha>0$. This includes classical examples of constant elasticity of substitution (CES), Cobb–Douglas, and Leontief utility functions.
The Nash welfare takes its name due to its connection with the bargaining problem introduced by Nash (1950). Nash studied general cooperative bargaining problems where the set of possible outcomes is a convex set $K\subseteq\mathbb{R}^{n}$ with a disagreement point $d\in K$ and showed that maximizing the geometric mean of the excess utilities above the disagreement utility amounts to the only possible bargaining outcome that satisfies certain natural requirements. Taking $K$ as the set of feasible allocations and $d=\mathbf{0}$, the Nash bargaining solution is precisely the Nash welfare maximizing allocation in this setting Vazirani (2012).
@@
Weak gross substitutes (WGS) utility functions are another broad class where a CE can be computed in polynomial time. Several fast algorithms have been designed for computing equilibria with WGS utilities Bei et al. (2019a), Codenotti et al. (2005b, c), Garg et al. (2023a). While the WGS class includes many homogeneous utilities, there are also non-homogeneous examples for which the Eisenberg–Gale solution may not correspond to a CE.
WGS and homogeneous utility functions cover most known cases where a CE can be computed in polynomial time. Already for the seemingly simple case of separable piecewise concave utilities, computing a CE is PPAD-hard Chen and Teng (2009). In fact, it is PPAD-hard even for bounded linear utilities, when there is an upper bound on the utility from each good, i.e., $u_{i}(x)=\sum_{j\in{\mathcal{M}}}\min\{v_{ij}x_{j},\ell_{ij}\}$ (Bei et al., 2019b). Example A.1 shows that the set of CE can be non-convex and disconnected already for bounded linear utilities.
@@ 1.1 Our contributions
In this paper, we study the relationship between two fundamental concepts in the allocation of divisible goods: competitive equilibrium and maximum Nash welfare allocations. We exhibit a close relationship between these solution concepts well beyond the homogeneous case. First, we show a ‘price of anarchy’ type result Dubey (1986), Koutsoupias and Papadimitriou (1999). One can interpret Nash welfare as a global measure of social utility. In contrast, a CE can be seen as an outcome reached by autonomous agents trying to improve their utilities individually. It turns out that, in the general setting of concave utilities, any CE recovers at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight (see Example A.4).
@@
We then study guarantees of the maximum Nash welfare solution in the context of CE. A major advantage of maximum Nash welfare is efficient computability: it can be arbitrarily well approximated for any concave utility functions. In contrast, computing CE becomes PPAD-hard already in very simple cases. In this context, our first simple result shows that the maximum Nash welfare yields an approximate CE for all concave utility functions. By an $\alpha$-demand-approximate CE, we mean allocations and prices such that every agent receives at least $1/\alpha$-fraction of the maximum possible utility (see Definition 2.8). As in the case of the Eisenberg–Gale program, one can interpret the optimal Lagrangian multipliers as prices. The result also implies approximate envy-freeness.
@@
For Fisher markets under concave utility functions, any Nash welfare maximizing allocation $\mathbf{y}$, together with the dual prices $\mathbf{q}$, form a 2-demand-approximate CE. This also implies that $\mathbf{y}$ is half envy-free when $b_{i}$’s are identical, i.e., $u_{i}(y_{i})\geq\tfrac{1}{2}u_{i}(y_{k})$, for all $i,k\in{\mathcal{A}}$.
Theorem 1.2 already shows that maximum Nash welfare provides remarkable fairness and efficiency guarantees for all concave utility functions. However, there can be multiple equilibria, giving vastly different utilities to an agent. In Example A.5, we show that there could be a gap of $\Omega(n)$ in the utility of some agent between two different competitive equilibria. Hence, an agent could be much better off in some CE than in the maximum Nash welfare solution.
As the main contribution of the paper, we show that maximum Nash welfare provides surprisingly strong guarantees for a broad class of utility functions. For this, we explore the Gale demand correspondence Garg et al. (2023a), Nesterov and Shikhman (2018). For an agent $i\in{\mathcal{A}}$ with budget $b_{i}$ and utility function $u_{i}(.)$ at prices $\mathbf{q}$, the Gale demand correspondence comprises of the allocations $y_{i}$ that maximize $b_{i}\log u_{i}(y_{i})-\langle\mathbf{q},y_{i}\rangle$. It follows from Lagrangian duality that in a Nash welfare maximizing solution, every agent receives an allocation from their Gale demand correspondence.
The reason behind the equivalence of CE and maximum Nash welfare for homogeneous utilities is that, for this class, the standard demand correspondence and the Gale demand correspondence coincide. Nevertheless, these two demands can behave very differently for non-homogeneous utility functions, e.g., the Gale demand can be non-monotonic in the prices, whereas the standard demand is always monotonic (see Example A.3).
@@
Competitive equilibrium becomes more subtle when the utility functions are satiable: there is an absolute maximum value of the utility that can be reached, regardless of the prices—bounded linear utilities mentioned above are such an example. To accommodate satiable utilities, we also introduce a stronger variant of the definition, called the $\Sigma$-Gale-substitutes property. We note that for non-satiable utilities, the Gale-substitutes and $\Sigma$-Gale-substitutes properties coincide. Figure 1 illustrates the containment relationships among various utility classes discussed in this paper, with those in which the Fisher market equilibrium can be efficiently computed highlighted in grey.
@@
Consider a Fisher market instance where all utility functions are $\Sigma$-Gale-substitutes. Then, for any competitive equilibrium $(\mathbf{x},\mathbf{p})$ and any Nash welfare maximizing allocation $\mathbf{y}$, $u_{i}(y_{i})\geq\frac{1}{2}u_{i}(x_{i})$ holds for every $i\in{\mathcal{A}}$.
As a counterpart to the ‘price of anarchy’ result in Theorem 1.1, one could call this a ‘price of authority’ result. The agents can either accept an allocation proposed by a central authority on the principle of maximizing Nash welfare or converge to a market solution (using some dynamics). By accepting the centrally proposed solution, each agent may be worse off by at most a factor 2 compared to the best possible equilibrium outcome. On the other hand, outcomes at different competitive equilibria can be very different, and the possible outcomes can be difficult to predict given the computational hardness of equilibria. Hence, for ($\Sigma$-)Gale-substitutes utility functions, the Nash welfare maximizing allocation provides a guaranteed good outcome to all agents simultaneously while also being efficiently computable. Thus, Theorem 1.3 establishes a strong relationship between CE and maximum Nash welfare beyond homogeneous functions and provides a natural alternative to the equilibrium selection problem Arrow and Hahn (1971) in this setting where multiple equilibria are possible.
We call a utility function $u(.)$ separable if $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. This class gives an important example of $\Sigma$-Gale-substitutes. Note that such utilities, in general, are neither homogeneous nor WGS. As mentioned above, competitive equilibrium computation is already PPAD-hard for the separable piecewise linear concave (SPLC) case when each $v_{j}$ is piecewise linear. This class has attracted significant attention on its own Anari et al. (2018), Chen et al. (2009, 2013), Garg et al. (2015), Vazirani and Yannakakis (2011).
@@
We consider two further examples of piecewise linear Gale-substitutes utilities:
  • Leontief-free (LF) utilities Filos-Ratsikas et al. (2024), Garg et al. (2014), Goyal et al. (2023): Consider a set $K$ of segments, and let $u(x)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\max\sum_{k\in K}u_{k}$: $u_{k}=\sum_{j\in{\mathcal{M}}}a_{jk}z_{jk},\ \ u_{k}\leq\ell_{k};\ \forall k\in K$; $\sum_{k\in K}z_{jk}\leq x_{j}\ \forall j\in{\mathcal{M}}$; $\mathbf{z}\geq 0$. That is, an LF function is given by a set of linear functions $u_{k}$’s with a maximum utility limit $\ell_{k}$ on each $k\in K$. LF captures interesting non-separable PLC functions, and includes SPLC as a special case where each segment $k\in K$ has exactly one $j$ with $a_{jk}>0$.
  • Generalized network utilities: We introduce a new and broad class of utility functions defined by generalized flow instances. Here, the goods correspond to source nodes in a flow network with supply $x_{j}$, $j\in{\mathcal{M}}$. Each arc has a capacity and a gain factor, and the utility $u(x)$ is the maximum flow value that can reach a designated sink node $t$ from supplies $x$. The precise definitions are given in Section 2.1. LF can be formulated as a special case of a network with two layers, the middle layer between the sources and the sink corresponding to the segments $K$.
@@
Generalized flows is a classical network flow model that can model commodity transportations with losses, as well as conversions between different types of entities such as in financial networks; see Chapter 15 in Ahuja et al. (1993) for a description and various applications.
Theorem 1.3 already shows that in Fisher markets with non-satiable generalized network utilities, Nash welfare maximizing allocations give a good approximation to utilities at competitive equilibrium. However, satiable generalized network utilities may not satisfy the stronger $\Sigma$-Gale-substitutes property required in Theorem 1.3. Example A.10 shows that the theorem does not hold if only assuming the Gale substitutes property. For satiable utility functions, it is natural to look at thrifty (strong) equilibria, which are guaranteed to exist Garg et al. (2022b). In such equilibria, each agent receives an optimal bundle at the cheapest possible price (see Definition 2.11). An appealing property of this concept is that thrifty equilibria are Pareto efficient, whereas this may not hold for (weak) competitive equilibrium. We also show that the statement of Theorem 1.3 remains true for generalized network utilities and for thrifty equilibria.
@@
Consider a Fisher market instance where all utility functions are generalized network utilities. Then, for any thrifty competitive equilibrium $(\mathbf{x},\mathbf{p})$ and any Nash welfare maximizing allocation $\mathbf{y}$, $u_{i}(y_{i})\geq\frac{1}{2}u_{i}(x_{i})$ holds for every $i\in{\mathcal{A}}$.
Finally, we note that the statement of Theorem 1.3 does not hold for arbitrary concave utilities. In Example A.8, we show that without the Gale-substitutes property, agents may get much higher utilities in CE than in the maximum Nash welfare allocations.
@@
We note that no polynomial time approximation scheme (PTAS) can exist that, for some $\varepsilon>0$, provides any agent at least $(1-\varepsilon)$ times their maximum CE utility. This is due to the fact that, as shown in Example A.1, multiple thrifty equilibria may exist, each providing different utilities to the agents. The existence of an allocation that $\varepsilon$-approximates the best CE utility for each agent would, by compactness, imply the existence of an allocation where every agent achieves their best CE utility. However, this would contradict the fact that thrifty equilibria are Pareto efficient.

### v3

@@ Abstract
We study the relationship between two central concepts in the allocation of divisible goods: competitive equilibrium (CE) and allocations that maximize Nash welfare, i.e., allocations where the weighted geometric mean of the utilities is maximal. When agents have homogeneous concave utility functions, these concepts coincide: the classical Eisenberg–Gale convex program that maximizes Nash welfare over feasible allocations yields a competitive equilibrium. However, they diverge for non–homogeneous utilities. From a computational perspective, maximizing Nash welfare amounts to solving a convex program for any concave utility functions, whereas computing CE becomes PPAD-hard already for separable piecewise linear concave (SPLC) utilities.
We introduce the concept of Gale-substitute utility functions, an analogue of the weak gross substitutes (WGS) property for the so-called Gale demand system. For Gale-substitutes utilities, we show that any allocation maximizing Nash welfare provides an approximate-CE with surprisingly strong guarantees, where every agent gets at least half the maximum utility they can get at any CE, and is approximately envy-free. Gale-substitutes include utility functions where computing CE is PPAD hard, such as all separable concave utilities and the previously studied non-separable class of Leontief-free utilities. We introduce a broad new class of utility functions called generalized network utilities based on the generalized flow model. This class includes SPLC and Leontief-free utilities, and we show that all such utilities are Gale-substitutes.
Conversely, although some agents may get much higher utility at a Nash welfare maximizing allocation than at a CE, we show a ‘price of anarchy’ type result: for general concave utilities, every CE achieves at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight.
@@ 1 Introduction
We investigate a setting where a set $\mathcal{M}$} of divisible goods needs to be fairly allocated among a set $\mathcal{A}$} of $n$ agents in a cardinal utility framework. Each agent $i\in{\mathcal{A}}$ has preferences given by a utility function $u_{i}(.)\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$. We make the standard assumptions that $u_{i}(.)$’s are monotone non-decreasing, concave, and satisfy $u_{i}(\mathbf{0})=0$. [^1] All concepts discussed in the Introduction will be formally defined in Section 2. Two classical allocation concepts are Nash welfare and competitive equilibrium from equal incomes (CEEI). In Nash welfare, we select an allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ that maximizes $\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{x}_{i})$. In CEEI, each agent is initially given an equal share of all goods and allowed to trade with others; the resulting allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ corresponds to a competitive equilibrium (CE) outcome.
Both these concepts have desirable properties that have made them widely popular in fair division research. Nash welfare derives its name from its roots in cooperative bargaining, introduced by Nash (1950). In this model, a convex set $K\subseteq\mathbb{R}^{n}$ of feasible outcomes is given, along with a disagreement point $\mathbf{d}\in K$ and concave utility functions $u_{i}\,:\,K\to\mathbb{R}_{\geq 0}$. Nash showed that the unique bargaining solution satisfying natural axioms—Pareto optimality, invariance under affine transformations, symmetry, and independence of irrelevant alternatives—is the one that maximizes the geometric mean of the agents’ gains over the disagreement point. Taking $K$ as the set of feasible allocations and $\mathbf{d}=\mathbf{0}$, the Nash bargaining solution exactly corresponds to the Nash welfare maximizing allocation in this setting (see also Vazirani (2012)).
CEEI, introduced by Varian (1974), also satisfies desirable properties such as Pareto optimality, scale invariance, and symmetry. In addition, it guarantees envy-freeness, which may not be the case for Nash welfare. For further discussion and references on these concepts, see Moulin (2003, Chapters 3 and 7).
Nash welfare maximization and CEEI remarkably coincide for degree one positively homogeneous utilities, i.e., if $u(\alpha\mathbf{x})=\alpha u(\mathbf{x})$ for any $\alpha>0$. This was first shown by Eisenberg and Gale (1959) for linear utilities (i.e., $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}x_{j}$), and later extended by Eisenberg (1961) to the homogeneous case that includes classical examples of constant elasticity of substitution (CES), Cobb–Douglas, and Leontief utility functions.
In general, Nash welfare maximization and CEEI can lead to different outcomes. Moreover, while the Nash welfare solution yields a unique utility profile for the agents, CEEI may admit multiple equilibria with significantly different utilities across agents. This raises a nontrivial equilibrium selection problem Arrow and Hahn (1971), Harsanyi and Selten (1998).
The problem of efficiently computing equilibria has been an important topic in economics, optimization, and computer science over the past decades. The existence of CEEI follows from classical results in general equilibrium theory by Arrow and Debreu (1954) and by McKenzie (1954), based on fixed point theorems. While these results apply to more general exchange market models, they do not yield efficient computational methods. In fact, the computational complexity of finding equilibria is generally negative: even for the seemingly simple case of separable piecewise concave (SPLC) utilities, computing a CE falls into the class of PPAD-hard Chen and Teng (2009) problems, suggesting that an efficient algorithm is unlikely to exist. The hardness persists even under bounded linear utilities, where the utility from each good is capped, i.e., $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}\min\{v_{j}x_{j},\ell_{j}\}$ (Bei et al., 2019b). Examples E.1 and E.2 show that the set of CE can be non-convex and disconnected even under such preferences. In contrast, Nash welfare maximization corresponds to solving a convex program and can therefore be computed efficiently for any concave valuations.
@@
Besides homogeneous utilities, another important class where CEEI can be computed in polynomial time is weak gross substitutes (WGS) utility functions. [^2] A utility function satisfies the weak gross substitutes property when raising the price of one good does not decrease the demand for any other good. Several fast algorithms have been designed for computing equilibria with WGS utilities Bei et al. (2019a), Codenotti et al. (2005b, c), Garg et al. (2023a). While the WGS class includes many homogeneous utilities, there are also non-homogeneous examples where the Eisenberg–Gale solution may not correspond to a CE. Together, WGS and homogeneous utility functions cover the majority of known settings in which a CE can be computed efficiently.
In this paper, we study the relationship between two fundamental concepts in the allocation of divisible goods: competitive equilibrium and maximum Nash welfare allocations. We reveal a strong connection between these solution concepts that extend well beyond the classical homogeneous case.
While the classical fair division concepts are CEEI and Nash welfare maximizing allocations, our results apply more broadly to the Fisher market model and weighted Nash welfare maximization. In the Fisher model, a fixed supply of goods $\mathcal{M}$} is to be allocated among agents $\mathcal{A}$}, where each agent $i\in{\mathcal{A}}$ has cardinal preferences given by a utility function $u_{i}(.)$. Each agent also has a fixed budget $b_{i}>0$. At competitive equilibrium (CE), prices $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ of goods and fractional allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ satisfy two conditions: (i) each agent $i$ gets an optimal bundle $\mathbf{x}_{i}$ that maximizes $u_{i}(\mathbf{x}_{i})$ subject to the budget constraint $\langle\mathbf{p},\mathbf{x}_{i}\rangle\leq b_{i}$, and (ii) the market clears, meaning demand meets supply for every good. The CEEI setting corresponds to the special case where all budgets are equal, i.e., $b_{i}=1$ for all $i\in{\mathcal{A}}$. Similarly, for given weights $b_{i}>0$, the weighted Nash welfare maximizing allocation is a feasible allocation of goods that maximizes $\prod_{i}u_{i}(\mathbf{x}_{i})^{b_{i}}$, or equivalently, $\sum_{i}b_{i}\log{u_{i}(\mathbf{x}_{i})}$. Classical results in Eisenberg and Gale (1959), Eisenberg (1961) establish that these allocations coincide with CE outcomes for degree one positively homogeneous utilities and arbitrary budgets $b_{i}>0$. Henceforth, by Nash welfare maximizing allocation we refer to this weighted concept.
@@ 1.1 Our contributions
Our first result is a ‘price of anarchy’ type bound Dubey (1986), Koutsoupias and Papadimitriou (1999). One can interpret Nash welfare as a global measure of social utility. In contrast, a CE can be seen as an outcome reached by autonomous agents trying to improve their utilities individually. It turns out that, in the general setting of concave utilities, any CE recovers at least $(1/\mathrm{e})^{1/\mathrm{e}}>0.69$ fraction of the maximum Nash welfare, and this factor is tight (see Example 3.2).
@@
We then study guarantees of the maximum Nash welfare solution in the context of CE. As noted above, a major advantage of maximum Nash welfare is efficient computability: it can be arbitrarily well approximated for any concave utility functions, whereas computing CE becomes PPAD-hard already in very simple cases. In this context, our first simple result shows that the maximum Nash welfare yields an approximate CE for all concave utility functions. As in the case of the Eisenberg–Gale program, one can interpret the optimal Lagrangian multipliers as prices. By an $\alpha$-demand-approximate CE, we mean allocations and prices such that every agent receives at least $1/\alpha$-fraction of the maximum possible utility at those prices (see Definition 2.8). The result also implies approximate envy-freeness.
@@
For Fisher markets under concave utility functions, any Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$, together with the dual prices $\mathbf{q}$, form a 2-demand-approximate CE. This also implies that $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is half envy-free when $b_{i}$’s are identical, i.e., $u_{i}(\mathbf{y}_{i})\geq\tfrac{1}{2}u_{i}(\mathbf{y}_{k})$, for all $i,k\in{\mathcal{A}}$.
Theorem 1.2 already shows that maximum Nash welfare provides remarkable fairness and efficiency guarantees for all concave utility functions. However, there can be multiple equilibria, giving vastly different utilities to an agent. In Example 4.4, we show that there could be a gap of $\Omega(n)$ in the utility of some agent between two different competitive equilibria. Hence, some agent could be much better off in some CE than in the maximum Nash welfare solution.
As the main contribution of the paper, we show that maximum Nash welfare provides surprisingly strong guarantees for a broad class of utility functions. For this, we explore the Gale demand correspondence Garg et al. (2023a), Nesterov and Shikhman (2018). For an agent $i\in{\mathcal{A}}$ with budget $b_{i}$ and utility function $u_{i}(.)$ at prices $\mathbf{q}$, the Gale demand correspondence [^3] In contrast, the standard demand correspondence comprises of the allocations $\mathbf{x}_{i}$ that maximize $\{u_{i}(\mathbf{x}_{i})\ |\ \langle\mathbf{q},\mathbf{x}_{i}\rangle\leq b_{i}\,\}.$ comprises of the allocations $\mathbf{y}_{i}$ that maximize
$b_{i}\log u_{i}(\mathbf{y}_{i})-\langle\mathbf{q},\mathbf{y}_{i}\rangle\,.$
It follows from Lagrangian duality that in a Nash welfare maximizing solution, every agent receives an allocation from their Gale demand correspondence.
The reason behind the equivalence of CE and maximum Nash welfare for homogeneous utilities is that, for this class, the standard demand correspondence and the Gale demand correspondence coincide. Nevertheless, these two demands can behave very differently for non-homogeneous utility functions, e.g., the utility value from the standard demand is always monotone non-increasing, whereas the utility under Gale demand can be non-monotonic (see Example E.3).
@@
Competitive equilibrium becomes more subtle when the utility functions are satiable: there is an absolute maximum value of the utility that can be reached, regardless of the prices—bounded linear utilities mentioned above are such an example. To accommodate satiable utilities, we also introduce a stronger variant of the definition, called the $\Sigma$-Gale-substitutes property. We note that for non-satiable utilities, the Gale-substitutes and $\Sigma$-Gale-substitutes properties coincide. Figure 1 illustrates the containment relationships among various utility classes discussed in this paper.
@@
Consider a Fisher market instance where all utility functions are $\Sigma$-Gale-substitutes. Then, for any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ and any Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$, $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}u_{i}(\mathbf{x}_{i})$ holds for every $i\in{\mathcal{A}}$.
As a counterpart to the ‘price of anarchy’ result in Theorem 1.1, one could call this a ‘price of authority’ result. The agents can either accept an allocation proposed by a central authority on the principle of maximizing Nash welfare or converge to a market solution (using some dynamics). By accepting the centrally proposed solution, each agent may be worse off by at most a factor 2 compared to the best possible equilibrium outcome. On the other hand, outcomes at different competitive equilibria can be very different, and the possible outcomes can be difficult to predict given the computational hardness of equilibria. Hence, for ($\Sigma$-)Gale-substitutes utility functions, the Nash welfare maximizing allocation provides a guaranteed good outcome to all agents simultaneously while also being efficiently computable. Thus, Theorem 1.3 establishes a strong relationship between CE and maximum Nash welfare beyond homogeneous functions and provides a natural alternative to the equilibrium selection problem Arrow and Hahn (1971), Harsanyi and Selten (1998) in this setting where multiple equilibria are possible.
We note that the statement of Theorem 1.3 does not hold for arbitrary concave utilities. In Example E.5, we show that without the Gale-substitutes property, agents may get much higher utilities in CE than in the maximum Nash welfare allocations.
Examples of Gale-substitute utilities
It turns out that ($\Sigma$-)Gale-substitutes capture a wide range of natural utility functions. We call a utility function $u(.)$ separable if $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. This class gives an important example of $\Sigma$-Gale-substitutes. Note that such utilities, in general, are neither homogeneous nor WGS. As mentioned above, competitive equilibrium computation is already PPAD-hard for the separable piecewise linear concave (SPLC) case when each $v_{j}$ is piecewise linear. This class has attracted significant attention on its own Anari et al. (2018), Chen et al. (2009, 2013), Garg et al. (2015), Vazirani and Yannakakis (2011).
@@
We consider two further examples of non-separable, piecewise linear Gale-substitutes utilities:
  • Leontief-free (LF) utilities Filos-Ratsikas et al. (2024), Garg et al. (2014), Goyal et al. (2023): Consider a set $K$ of segments, and let $u(\mathbf{x})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\max\sum_{k\in K}u_{k}$: $u_{k}=\sum_{j\in{\mathcal{M}}}a_{jk}z_{jk},\ \ u_{k}\leq\ell_{k};\ \forall k\in K$; $\sum_{k\in K}z_{jk}\leq x_{j}\ \forall j\in{\mathcal{M}}$; $\mathbf{z}\geq 0$. That is, an LF function is given by a set of linear functions $u_{k}$, with a maximum utility limit $\ell_{k}$ on each $k\in K$. LF captures interesting non-separable PLC functions, and includes SPLC as a special case where each segment $k\in K$ has exactly one $j$ with $a_{jk}>0$.
  • Generalized network utilities: We introduce a new and broad class of utility functions defined by generalized flow instances. Here, the goods correspond to source nodes in a flow network with supply $x_{j}$, $j\in{\mathcal{M}}$. Each arc has a capacity and a gain factor, and the utility $u(\mathbf{x})$ is the maximum flow value that can reach a designated sink node $t$ from supplies $\mathbf{x}$. This value can be computed in polynomial time. The precise definitions are given in Section 2.1. Additionally, LF can be formulated as a special case of a network with two layers, the middle layer between the sources and the sink corresponding to the segments $K$.
We note that generalized flows is a classical network flow model that can model commodity transportation with losses, as well as conversions between different types of entities such as in financial networks; see Ahuja et al. (1993, Chapter 15) for a description and various applications.
@@
Theorem 1.3 already shows that in Fisher markets with non-satiable generalized network utilities, Nash welfare maximizing allocations give a good approximation to utilities at competitive equilibrium. However, satiable generalized network utilities may not satisfy the stronger $\Sigma$-Gale-substitutes property required in Theorem 1.3. Example E.6 shows that the theorem does not hold if only assuming the Gale substitutes property. For satiable utility functions, it is natural to look at thrifty (strong) equilibria, a strengthening of CE, which are also guaranteed to exist, see, e.g., Garg et al. (2022b). In such equilibria, each agent receives an optimal bundle at the cheapest possible price (see Definition 2.11). An appealing property of this concept is that thrifty equilibria are Pareto efficient, whereas this may not hold for (weak) competitive equilibrium. We also show that the statement of Theorem 1.3 remains true for generalized network utilities and for thrifty equilibria.
@@
Consider a Fisher market instance where all utility functions are generalized network utilities. Then, for any thrifty competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ and any Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$, $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}u_{i}(\mathbf{x}_{i})$ holds for every $i\in{\mathcal{A}}$.
@@
We note that no algorithm can guarantee agents at least $(1-\varepsilon)$ times their maximum CE utility for any $\varepsilon>0$. This is due to the fact that, as shown in Example E.1, multiple thrifty equilibria may exist, each providing different utilities to the agents. The existence of an allocation that $\varepsilon$-approximates the best CE utility for each agent would, by compactness, imply the existence of an allocation where every agent achieves their best CE utility. However, this would contradict the fact that thrifty equilibria are Pareto efficient.